OpenStudy (anonymous):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. My equation is f(x)=-2x^2+40x-182 in standard form but im lost, I dont know what to do?

OpenStudy (anonymous):

Wouldn't i have to just plug in the numbers for x?

OpenStudy (jadeishere):

lol is this algebra 1?

OpenStudy (anonymous):

Algebra 2

OpenStudy (anonymous):

I believe you need to use the slope formula...

OpenStudy (anonymous):

average rate of change is f(b)−f(a)b−a so substitute the x values to get a corresponding f(x) then substitute into the formula

OpenStudy (anonymous):

The x values would be -134 and 0 correct?

OpenStudy (anonymous):

Yes. I do believe that is correct. :)

OpenStudy (anonymous):

So what would i do now?

OpenStudy (anonymous):

0-(-134)?

OpenStudy (anonymous):

so all you need to do for the average rate of change is find the slope of the interval joining the 2 points. I think..

OpenStudy (anonymous):

Is there an equation for that? lol

OpenStudy (anonymous):

Can I give you a example of one? Lol

OpenStudy (anonymous):

F(7)=3(7−3)2+4=52 and for x = 4 F(4)=3(4−3)2+4=7 so the average rate of change is 52−7 7−4=

OpenStudy (anonymous):

Oh okay. Thank you! :)

OpenStudy (anonymous):

Your welcome!